OpenStudy (anonymous):

-x-y=1 Graph the linear equation Neeed help ASAP

OpenStudy (anonymous):

it is easier to graph these if they are in slope intercept form, y = mx + b -x -y = 1 +x on both sides -y = x+1 to get positive y, multiply by -1 on both sides: y = -x-1 slope is -1 and it intercepts the y-axis at -1, do you know how to graph it now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hey does this help?

OpenStudy (anonymous):